OpenStudy (anonymous):

Find the slope of the line that passes through the points (1,4) and (4,7). Answer -3/2 1 3/2 2

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You need to know how to calculate the slope. The y-value changes and the x-value changes from one point to the other. What are these values?

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=(7-4)/(4-1) = 9

OpenStudy (anonymous):

oops, m=1

OpenStudy (anonymous):

lol sorry :)

OpenStudy (anonymous):

I was gonna say @privetek (: haha it cool.